OpenStudy (anonymous):

Three times the sum of 4 and a number y squared is the same as the difference of 14and the number y

OpenStudy (anonymous):

3(4+y^2)=14-y 12 - 3y^2 = 14 - y 3y^2 - y +2 = 0 y = 0.167 + 0.799 i or y = 0.167 - 0.799 i

OpenStudy (anonymous):

tutor your supposed to simplify